OpenStudy (anonymous):

is this right 9. Solve x/14 + 3 = -1 for x. (Points : 4) -52 -56 -60 -64<

OpenStudy (anonymous):

No.

jimthompson5910 (jim_thompson5910):

are you saying that the answer is -64 ?

OpenStudy (anonymous):

@jim_thompson5910 ya

jimthompson5910 (jim_thompson5910):

if so, then let's check it x/14 + 3 = -1 -64/14 + 3 = -1 ... plug in x = -64 -4.57142857 + 3 = -1 -1.57142857 = -1 ... which is FALSE So x = -64 isn't the answer

OpenStudy (anonymous):

-56 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's check that x/14 + 3 = -1 -56/14 + 3 = -1 ... plug in x = -56 -4 + 3 = -1 -1 = -1 ... which is TRUE So x = -56 is the answer